Alyeska Services Company, a division of a major oil company, provides various services to the operators of the North Slope oil field in Alaska. Data concerning the most recent year appear below:


Sales $17,700,000
Net operating income $4,900,000
Average operating assets $35,300,000

Requirement 1:
Compute the margin for Alyeska Services Company. ((Round your answer to 2 decimal places.)

Requirement 2:
Compute the turnover for Alyeska Services Company. (Round your answer to 2 decimal places.)

Requirement 3:
Compute the return on investment (ROI) for Alyeska Services Company. (Round your intermediate calculations and final answers to two decimal places)
Answer #1

1.

Margin = Net operating income /sales

= 4,900,000 / 17,700,000

= 27.68%

2.

Turnover = Sales / Average operating assets

= 17,700,000 / 35,300,000

= 0.50

3.

ROI = margin * turnover

= 27.68*0.5

= 13.84%

I have rounded some numbers. Comment if you face any issues